OpenStudy (anonymous):

7t^9 Divided by t^6 . Help me out please?

OpenStudy (anonymous):

you have to use this rule of exponenets : a^m/a^n=a^m-n

jimthompson5910 (jim_thompson5910):

\[\large \frac{7t^9}{t^6}\] \[\large 7t^{9-6}\] \[\large 7t^3\] So \[\large \frac{7t^9}{t^6}=7t^3\]

OpenStudy (anonymous):

Or you can use the multiplication law

OpenStudy (anonymous):

in that case you know that 1/t^6 = t^-6 so you get 7t^9*t^-6 (when multiplying two of the same base terms withexponents you add so it's be \[7t^(9+-6) = 7t^3\]

jimthompson5910 (jim_thompson5910):

type \ [ 7^{-6} = \frac{1}{7^6} \ ] to get \[\large 7^{-6} = \frac{1}{7^6} \]

jimthompson5910 (jim_thompson5910):

oh and remove all spaces from the code above

OpenStudy (anonymous):

gotcha... why don't they have a fraction button in the menu >_<

jimthompson5910 (jim_thompson5910):

everything you'll need can be found in the equation editor by clicking on the "equation" button below

jimthompson5910 (jim_thompson5910):

good question, didn't realize it wasn't there

OpenStudy (anonymous):

haha yeah it'd be useful

jimthompson5910 (jim_thompson5910):

for more functions, look up LaTex since this is the language it's using